OpenStudy (anonymous):

Bobbi saved $1,000 from her summer job, she puts it into an account with a 6.9% interest compounded annually. How much will she have in her account after five years?

OpenStudy (campbell_st):

do you know the compound interest formula...?

OpenStudy (anonymous):

is it f(t)=a(1+b)^t?

OpenStudy (campbell_st):

ok... so a = $1000 b = 6.9% as a decimal t = 5 years... just substitute and calculate... and make sure you have b as a decimal when substituting

OpenStudy (anonymous):

Is the answer about $1396?

ganeshie8 (ganeshie8):

1396 is \(\large \color{red}{\checkmark}\) http://www.wolframalpha.com/input/?i=1000%281+%2B+0.069%29%5E5

ganeshie8 (ganeshie8):

good job !

OpenStudy (wolf1728):

Here's a good site for calculating compound interest: http://1728.org/compint.htm

OpenStudy (anonymous):

Thank you! Oh and since you are here can you help me out with similar ones that go along with this question? 2. What if the interest was 6.87% compounded quarterly (4 times a year)

ganeshie8 (ganeshie8):

use below formula : \(\large A = P(1 + \frac{r}{n})^{n\times t}\)

OpenStudy (anonymous):

it says math processing error...

ganeshie8 (ganeshie8):

A = P(1 + r/n)^(nt)

ganeshie8 (ganeshie8):

n = 4 here

OpenStudy (wolf1728):

try mine - it even has quarterly compounding

OpenStudy (anonymous):

Ok thank you. So now what if the interest was 6.85% compounded monthly?

ganeshie8 (ganeshie8):

how many months are there in a year ?

OpenStudy (anonymous):

12

ganeshie8 (ganeshie8):

so, n = 12

ganeshie8 (ganeshie8):

plugit in the formula : A = P(1 + r/n)^(nt) = 1000( 1 + 0.0685/12)^(12*5)

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

I got $1410

OpenStudy (anonymous):

oh wait hold on

ganeshie8 (ganeshie8):

use wolfram as ur calculator : http://www.wolframalpha.com/input/?i=1000%281+%2B+0.0685%2F12%29%5E%285*12%29

OpenStudy (anonymous):

ok i got $27961

OpenStudy (anonymous):

ugh I put the percent in wrong

ganeshie8 (ganeshie8):

try again, u should get 1407.09

ganeshie8 (ganeshie8):

1000 dollars wont become 27961 dollars in just 5 years !!! xD

OpenStudy (anonymous):

yup got it!! Thank you so much. Okay I only have 2 more. What about 6.8$ interest compounded daily (365 times a year--leap year not included)

ganeshie8 (ganeshie8):

plugin n = 365

OpenStudy (anonymous):

*6.8%

ganeshie8 (ganeshie8):

use same formula

OpenStudy (anonymous):

ok i got $1404

OpenStudy (wolf1728):

I got 1,404.90

ganeshie8 (ganeshie8):

always trust wolfram :)

ganeshie8 (ganeshie8):

1404.9 ~ 1405

OpenStudy (anonymous):

Yay okay! One more!! What if the interest was 6.7% compounded continuously?

ganeshie8 (ganeshie8):

now the fun part :)

ganeshie8 (ganeshie8):

\(e\) kicks in

OpenStudy (anonymous):

it says math processing error

ganeshie8 (ganeshie8):

"e" kicks in

OpenStudy (anonymous):

ohhh okay. So what do I put in wolfram lol

ganeshie8 (ganeshie8):

whenever u see "continuous", it got to do with "e"

ganeshie8 (ganeshie8):

A = P*e^(r*t)

ganeshie8 (ganeshie8):

P = initial investment r = interest rate t = # of years

ganeshie8 (ganeshie8):

plugin

OpenStudy (anonymous):

i got 3.53887... *10^17

OpenStudy (anonymous):

????

ganeshie8 (ganeshie8):

wolfram says 1397.94 http://www.wolframalpha.com/input/?i=1000%28e%29%5E%280.067*5%29

OpenStudy (anonymous):

yup thats what it said.. YAY Thank you so much!!!!!

ganeshie8 (ganeshie8):

np.. u wlc :)